40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2799     Entry time: Mon Aug 23 14:35:36 2021
Author: shruti 
Type: DailyProgress 
Category: PSOMA 
Subject: re-aligning 

[shruti, aaron]

After plugging back in all the cables, I was struggling to get the cavity locked again or seeing the transmitted beam on the monitor.

Today, following Aaron's suggestion, I removed the connection from the PDH mixer IF port to the oscilloscope and also the power splitter I had added earlier (now sending the IF output directly to the LB input only with the 10 dB attenuator). The cavity didn't lock right away but changing the gain to 6 and re-adjusting the input and output offset got the cavity locked again.

When we tried to get it on the monitor, we saw that the mode it locked to was a higher order mode misaligned in the vertical direction. We first aligned the beam vertically until we started seeing lower order modes and finally the TEM00 mode.

I adjusted the lenses a little to tweak the mode-matching after aligning the beam into the cavity for maximum transmission when locked. Calculating the mode-matching using the reflected beam when locked and unlocked shows that it is only 60% though.

 

ELOG V3.1.3-